Over the years, O'Brien Corporation's stockholders have provided $20,000,000 of capital, when they purchased new issues of stock and allowed management to retain some of the firm's earnings. The firm now has 1,000,000 shares of common stock outstanding, and it sells at a price of $32.00 per share. How much value has O'Brien's management added to stockholder wealth over the years, i.e., what is O'Brien's MVA?
Answer #1

The market value of equity = Stock price × Number of shares =1000000*32 = 32,000,000

MVA = Market value of equity − Book value of equity =32,000,000-20,000,000 =12,000,000